OpenStudy (anonymous):

The perimeter of a rectangle is 132m the length is 6m more than twice the width find the dimensions. What is the length? What is the width?

OpenStudy (anonymous):

P=(l+W)*2 l=6+2W 132=(6+2W)+W)*2 132/2 =6+3W 66 -6 =3W 60/3 = W 20 = W L = 6+(2*20) L = 46

OpenStudy (anonymous):

click good answeer if I deserve

OpenStudy (anonymous):

What about the width?

OpenStudy (anonymous):

width = 20 length=46